# Submodular Maximization over a Matroid $k$-Intersection: Multiplicative Improvement over Greedy

### v1

@@ Appendix A Polynomial Time Implementation of Algorithm 1
For every iteration $i$ in which Algorithm 3 produces a set $A_{i}$, we have $f(e\mid A_{\leq i})<m_{i}$ for all $e\in E\setminus A_{\leq i}$ that obey $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration ii.
Our next objective is to show that Algorithm 3 correctly simulates Algorithm 1. We assume in the proof that the two algorithms use the same method for finding the next $(m_{i},\varepsilon)$-improvement in each iteration of the inner loop, and we fix the values of any random bits used by this method (if it is randomized), as well as the value of $\alpha$. Once these values are fixed, showing that Algorithm 3 correctly simulates Algorithm 1 boils down to proving that they both return the same output set. Intuitively, to prove this, we need to show that the “fast forward” operation done by Line 3 of Algorithm 3 works as expected by the intuition given at the beginning of this appendix. The next lemma is the first step towards this goal. It proves the simpler claim that the fast forward operation always increases ii. Notice that this lemma is necessary for guaranteeing that each iteration of Algorithm 3 has a unique index ii.
@@
Every execution of Line 3 of Algorithm 3 strictly increases the value of ii. Moreover, after $i$ is increased, we have $m_{i}\leq w_{i^{\prime}}<m_{i-1}=2m_{i}$
@@
By Lemma A.4, $i$ never reaches any value larger than $L$. Together with the definition of $C$, this implies that the number of iterations Algorithm 3 performs is $|C|$. By Lemma A.3, each value $i\in C$ corresponds to a non-empty set $A_{i}$ in Algorithm 1, and since these sets are disjoint by construction, $|C|\leq|E|$. Thus, Algorithm 3 makes at most $|E|$ iterations. Each such iteration takes at most $O(|E|)$ time (the time required to compute $w_{i}$ and update ii) plus the time required to search and apply the $(m_{i},\varepsilon)$ improvements. Thus, the running time of Algorithm 3 is $O(|E|^{2})$ plus the total time spent searching for and applying these improvements across all iterations.
Let us denote by NN the number of improvements found throughout the algorithm. Notice that in each iteration of Algorithm 3, the number of times that the algorithms looks for an improvement is larger by exactly 11 compared to the number of improvements found in this iteration. Thus, Algorithm 3 looks for improvements at most $M+|E|$ times. Looking for an improvement can be done in $O(|E|^{3})$ time by simply enumerating over all the possible options for a set $S\subseteq E\setminus(A_{i}\cup A_{\leq i^{\prime}})$ of size at most 22 and a set $N\subseteq A_{i}$ of size at most 11. Once an improvement is found, applying it (i.e., replacing $A_{i}$ with $(A_{i}\setminus N)\cup S$) can be done in $O(1)$ time, and thus, by the above discussion, the time complexity of Algorithm 3 is
@@
It remains to bound $M$. Consider the results of applying each of the three types of improvements in Definition 3.1 in some iteration ii.
@@
(recall that $x_{1}$ and $x_{2}$ are labels given by Definition 3.1 for the two elements of $S$, and $y$ is the single element of NN). Thus, if we consider $|A_{i}|$ and $f(A_{\leq i^{\prime}}\cup A_{i})$ as two potentials then no $(m_{i},\varepsilon)$-improvement ever decreases these potentials, and each such improvement either increases $|A_{i}|$ by 11 or increases $f(A_{\leq i^{\prime}}\cup A_{i})$ by at least $\varepsilon m_{i}$. Let us denote by $\hat{A}_{i}$ the final value of the set $A_{i}$ in iteration ii. Then, the potential $|A_{i}|$ is non-negative and upper bounded by $|\hat{A}_{i}|$, and therefore at most $|\hat{A}_{i}|$ $(m_{i},\varepsilon)$-improvements can increase it by 11. Similarly, by the submodularity of $f$, the potential $f(A_{\leq i^{\prime}}\cup A_{i})$ starts with a value of $f(A_{\leq i^{\prime}})$ and ends up with the value

### v2

@@ Appendix A Polynomial Time Implementation of Algorithm 1
For every iteration $i$ in which Algorithm 3 produces a set $A_{i}$, we have $f(e\mid A_{\leq i})<m_{i}$ for all $e\in E\setminus A_{\leq i}$ that obey $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration $i$.
Our next objective is to show that Algorithm 3 correctly simulates Algorithm 1. We assume in the proof that the two algorithms use the same method for finding the next $(m_{i},\varepsilon)$-improvement in each iteration of the inner loop, and we fix the values of any random bits used by this method (if it is randomized), as well as the value of $\alpha$. Once these values are fixed, showing that Algorithm 3 correctly simulates Algorithm 1 boils down to proving that they both return the same output set. Intuitively, to prove this, we need to show that the “fast forward” operation done by Line 3 of Algorithm 3 works as expected by the intuition given at the beginning of this appendix. The next lemma is the first step towards this goal. It proves the simpler claim that the fast forward operation always increases $i$. Notice that this lemma is necessary for guaranteeing that each iteration of Algorithm 3 has a unique index $i$.
@@
Every execution of Line 3 of Algorithm 3 strictly increases the value of $i$. Moreover, after $i$ is increased, we have $m_{i}\leq w_{i^{\prime}}<m_{i-1}=2m_{i}$
@@
By Lemma A.4, $i$ never reaches any value larger than $L$. Together with the definition of $C$, this implies that the number of iterations Algorithm 3 performs is $|C|$. By Lemma A.3, each value $i\in C$ corresponds to a non-empty set $A_{i}$ in Algorithm 1, and since these sets are disjoint by construction, $|C|\leq|E|$. Thus, Algorithm 3 makes at most $|E|$ iterations. Each such iteration takes at most $O(|E|)$ time (the time required to compute $w_{i}$ and update $i$) plus the time required to search and apply the $(m_{i},\varepsilon)$ improvements. Thus, the running time of Algorithm 3 is $O(|E|^{2})$ plus the total time spent searching for and applying these improvements across all iterations.
Let us denote by $N$ the number of improvements found throughout the algorithm. Notice that in each iteration of Algorithm 3, the number of times that the algorithms looks for an improvement is larger by exactly $1$ compared to the number of improvements found in this iteration. Thus, Algorithm 3 looks for improvements at most $M+|E|$ times. Looking for an improvement can be done in $O(|E|^{3})$ time by simply enumerating over all the possible options for a set $S\subseteq E\setminus(A_{i}\cup A_{\leq i^{\prime}})$ of size at most $2$ and a set $N\subseteq A_{i}$ of size at most $1$. Once an improvement is found, applying it (i.e., replacing $A_{i}$ with $(A_{i}\setminus N)\cup S$) can be done in $O(1)$ time, and thus, by the above discussion, the time complexity of Algorithm 3 is
@@
It remains to bound $M$. Consider the results of applying each of the three types of improvements in Definition 3.1 in some iteration $i$.
@@
(recall that $x_{1}$ and $x_{2}$ are labels given by Definition 3.1 for the two elements of $S$, and $y$ is the single element of $N$). Thus, if we consider $|A_{i}|$ and $f(A_{\leq i^{\prime}}\cup A_{i})$ as two potentials then no $(m_{i},\varepsilon)$-improvement ever decreases these potentials, and each such improvement either increases $|A_{i}|$ by $1$ or increases $f(A_{\leq i^{\prime}}\cup A_{i})$ by at least $\varepsilon m_{i}$. Let us denote by $\hat{A}_{i}$ the final value of the set $A_{i}$ in iteration $i$. Then, the potential $|A_{i}|$ is non-negative and upper bounded by $|\hat{A}_{i}|$, and therefore at most $|\hat{A}_{i}|$ $(m_{i},\varepsilon)$-improvements can increase it by $1$. Similarly, by the submodularity of $f$, the potential $f(A_{\leq i^{\prime}}\cup A_{i})$ starts with a value of $f(A_{\leq i^{\prime}})$ and ends up with the value